Curriculum Vitae of András RECSKI (December 2002)

Born: April 21, 1948 in Budapest, Hungary

Citizenship: Hungarian

Home address: H-1111 Budapest, Lágymányosi u. 16. V. 3. Hungary

Current position: Professor of mathematics, head of the department, Budapest University of Technology and Economics, Faculty of Electrical Engineering and Informatics, Department of Computer Science and Information Theory (since 1990)

Phone: (+36-)1-2097-341 (private), (+36-)1-4632-585 (office)

Fax (+36-)1-4633-157

E-mail recski@cs.bme.hu

Education: Diploma in mathematics (equivalent to Bachelor + Master), L. Eötvös University, Budapest, Hungary, 1971.

Degrees: Dr. Rer. Nat. (equivalent to Ph. D.), L. Eötvös University, Budapest, Hungary, 1973; Candidate of Math. Sci. and Doctor of Math. Sci. (postdoctoral degrees), Hungarian Academy of Sciences, Budapest, Hungary, 1977 and 1984, respectively.

Positions held: Member of research staff, Research Institute for Telecommunication, Budapest, Hungary, 1971-84; Lecturer (part time), L. Eötvös University, Budapest, Hungary, 1972-84; Professor of computer science, L. Eötvös University, Budapest, Hungary, since 1984 (part time since 1990); Professor of mathematics, Budapest University of Technology and Economics, Faculty of Electrical Engineering and Informatics, since 1990; External member, Mathematical Research Institute of the Hungarian Academy of Sciences, Budapest, Hungary, since 1983.

Visiting positions: Danmarks Tekniske Højskole, Lyngby, Denmark, 1975 and 1976; Middle East Technical University, Ankara, Turkey, 1977; Technische Hochschule Ilmenau, Germany, 1978 and 1980; University of Tokyo and Tokai University, Japan, 1978/79; University of Waterloo, Waterloo, Ontario, Canada 1984; Cornell University, Ithaca, New York, USA, 1985; University of Bonn, Germany, 1987-89 and 1998/99; Yale University, New Haven, Connecticut, USA, 1994-95.